OpenStudy (anonymous):

If Tuscon's average rainfall is 12 1/4 anf Yuma's is 2 3/5 how much more rain, on average, does Tuscon get then Yuma? Must be interger, proper fraction or mixed number. simplify. any help?

OpenStudy (anonymous):

(49/4)-(13/5), next find the common denominater , which is 20. So, we adjust each fraction: (245/20)-(52/20)=193/20

OpenStudy (anonymous):

oh yeah...helping cheaters :P

OpenStudy (anonymous):

ewwwww

OpenStudy (anonymous):

Which means on average: Tuscon has about (193/20)=9.56 inches of rain

OpenStudy (anonymous):

Sweetiee pieee is you sugar pumpkin

OpenStudy (anonymous):

uhahahaha)))

OpenStudy (anonymous):

Am I interrupting u?:P

OpenStudy (anonymous):

mixed fraction from: 9 and 13/20